Physical Chemistry

Second excitation potential for He+ ion is (1) 12.09 eV (2) 48.36 eV (3) 3.4 eV (4) 51 eV

Siddharth Girish Naidu
6 Years agoGrade

2 Answers

ROSHAN MUJEEB
6 Years ago
Dear Student
The formula for excitation potential between two states n1 and n2 is U= -13.6Z^2/n1^2-13.6Z^2/n2^2
Here Z=2, n1=1,n2=3. Here U=48.36
MANTHAN RAO
6 Years ago
Excitation potential is the amount of energy required in electron volt to excite the atom
from it's ground state to a higher state.
When there is transition from n = 1 to n = 2 we will have first excitation potential.
For hydrogen atom,
First excitation potential = -3.4-(-13.6) eV = 10.2 eV
When electron jumps from n = 1 level to n = 3 level, we will have second exciation
potential.
Second excitation potential =-1.5-(-13.6) = 12.09 eV.